Internet Tools Summary

John December (decemj@rpi.edu)

December 15, 1993

15 Dec 93; Release 1.35

Copyright \begin{matrix}c \\ ◯ \end{matrix}1993 by John December (decemj@rpi.edu). Permission to use for non-commercial, personal or educational purposes. This copyright and permission notice must appear in all copies. Permission is also granted to mention, cite, refer to, or describe this document in books, products, or online services (but not to reproduce in whole or in part without permission.)

DISCLAIMER: Provided “as is” without expressed or implied warranty. I make no representations about this document’s suitability, accuracy, or stability.

PURPOSE: to list tools available on the Internet that are used for network information retrieval (NIR), Computer-Mediated Communication (CMC), or other services. This is not meant to be a strict categorization or an exhaustive list, rather a reference catalog. I certainly would welcome comments and suggestions.

AUDIENCE: those getting started in understanding what you can do on the Internet in NIR and CMC; for those experienced, it collects and summarizes sources of information.

ASSUMPTION: you have access to and know how to use finger, ftp, gopher, http, telnet, email, or Usenet newsgroups.

REFERENCES: The references listed at the end of this document are very useful guides to these tools.

ACTION: The Action: field of the entries is described in the section at the end of this document, ACTION NOTATION. This is a notation that I’ve devised to try to characterize the tools further; I’m not quite satisfied with this notation, though, yet.

NOTES:

1. Respect your access privileges to these tools.

2. This information changes; additions/comments welcome.

3. For demo purposes, I have used Unix commands; certainly Unix is not the only operating system required for these tools and forums. Apologies to those who don’t have Unix.

4. On the classification (NIR v. CMC): some tools could be used for either, such as telnet and email. I’ve placed them in what I feel are ”principle use” categories, e.g., telnet is used mostly for NIR, and email is commonly used for CMC (although both have applications in the other category).

5. Many of these tools have applicability off the Internet. Usenet, for example, is not confined to the Internet, and Internet email (and thus LISTSERV files) can be exchanged with communication systems off the Internet (BITNET, fidonet, commercial services.) So this list is not Internet (only) tools, but tools that can be used on the Internet.

Contents

1 HOW TO USE THIS DOCUMENT 2 2 NETWORK INFORMATION RETRIEVAL (NIR) TOOLS 2 3 COMPUTER-MEDIATED COMMUNICATION (CMC) FORUMS 9 4 PROTOCOLS/STANDARDS 11 5 REFERENCES 12 6 ACTION NOTATION 13 7 DESCRIPTION OF ITEMS 14

1 HOW TO USE THIS DOCUMENT

This document is formatted for LaTEX(for conversion to Postscript.) See the readme file for information on different formats for this document.

The tables in this document present information about the Internet. Tables of networkaccessible information are in the following form:

NETWORK ACCESSIBLE INFORMATION

ITEM NAME AM

PARAMETERS PARAMETERS

Where • ITEM NAME = a short name describing the information or service.

• AM (ACCESS METHOD) = one of email, finger, ftp, gopher, http, phone, telnet, unix, Usenet news.

• PARAMETERS = further access information depending on ACCESS METHOD:

ACCESS METHOD AND PARAMETERS

ACCESS METHOD PARAMETERS PARAMETERS
email  address  “letter-body”
finger  address  port #
ftp  host  path-to-directory/file-name
gopher  host
http host path-to-directory/file-name
phone  phone #
telnet host port # ‘login-name’
unix  command
news  newsgroup-name

If any of the parameters are not on the line, they are not needed.

• All ftp access uses login-name ‘anonymous’ unless stated otherwise.

For information that is paper-based:

PAPER-BASED INFORMATION

Author year title

publisher or further information

INFORMATION ABOUT UPDATES OF THIS DOCUMENT

Internet-tools ftp ftp.rpi.edu pub/communications/internet-too ls.readme

2 NETWORK INFORMATION RETRIEVAL (NIR) TOOLS

Alex

Definition: Alex provides transparent read of remote files at anonymous FTP sites;

Action: U := F <- S[F]

Alex description ftp

alex.sp.cs.cmu.edu /usr0/anon/doc/NIR.Tool

Alex document ftp alex.sp.cs.cmu.edu /usr0/anon/doc/intro.ps

Archie

Definition: Archie locates files at anonymous FTP sites by filename search;

Action: U := M(L <= C <- S[F])

Archie description ftp

archie.ans.net pub/archie/doc/whatis.archie

Archie manual page ftp

archie.ans.net pub/archie/doc/archie.man.txt

Archie via telnet telnet archie.ans.net

‘archie’

Archie via WWW http web.nexor.co.uk

archie.html

Archie via Unix unix

archie -s STRING

Archie via email mail archie@archie.unl.edu

“help”

Astra

Definition: Astra retrieve documents from databases;

Action: U == M2 <- S[F] <- M1 <= U

Astra Help mail astradb@icnucevm.bitnet

“help”

Astra documentation mail astradb@icnucevm.bitnet “GET META DOCUMENT”

Biomix

Definition: Biomix presents network resources as locations on screen maps; Action: U :: G Biomix description ftp mailbase.ac.uk pub/lists/unite/files/biomix.tx t

Cello

Definition: Cello is a DOS-based Internet browser incorporating WWW, Gopher, FTP, Telnet, News; Action: U :: G Cello source ftp fatty.law.cornell.edu pub/LII/Cello/

Compass

Definition: Compass is a system for Internet access to resources and tools for CERFnet subscribers; Action: Compass description ftp mailbase.ac.uk pub/lists/unite/files/compass.t xt

Essence

Definition: Essence is a resource discovery system using indexes; Action: Essence description ftp

ftp.cs.colorado.edu pub/cs/distribs/essence/README

Essence document ftp ftp.cs.colorado.edu pub/cs/distribs/essence/Essence .txt.Z

Finger

Definition: Finger retrieves information about user registered on a host computer;

Action: U := M <-<= H[R]

Finger description unix

man finger

Finger demo finger

decemj@rpi.edu

Finger Protocol ftp

nic.merit.edu documents/rfc/rfc1288.txt

Finger via email mail

b.liddicott@ic.ac.uk  Subject:  #finger

USER@HOST.DOMAIN

Finger via WWW http

cs.indiana.edu finger/gateway?USER

Finger via telnet telnet HOST

79

FingerInfo ftp

csd4.csd.uwm.edu pub/fingerinfo

Who fingered? ftp

ftp.csie.nctu.edu.tw Upload/backfinger/

File Transfer Protocol (Ftp)

Definition: Ftp retrieves or puts copies of files at remote FTP sites;

Action: U == F <- H[F]; U => F -> H[F]

Ftp how to ftp ftp.sura.net pub/nic/network.service.guides/

how.to.ftp.guide

FTP ftp

nic.merit.edu documents/rfc/rfc0959.txt

Ftp via Unix unix ftp

Ftp via email mail ftpmail@decwrl.dec.com

“help”

Ftp via telnet telnet

grind.isca.uiowa.edu

File Compression ftp

ftp.cso.uiuc.edu doc/pcnet/compression

FTP setup ftp cert.org

pub/tech tips/anonymous ftp

Gopher

Definition: Gopher locates and retrieves resources using a graph of menus;

Action: U ::  I(L <- C <= S[F]); U := M <- S[F]

Gopher guide ftp boombox.micro.umn.edu pub/gopher/docs/GopherGuide Jun 15b.txt

Gopher demo gopher

gopher.micro.umn.edu

Gopher via telnet telnet consultant.micro.umn.edu

‘gopher’

Gopher via email mail gophermail@ncc.go.jp

“help”

Gopher source ftp

boombox.micro.umn.edu pub/gopher/00README

Gopher documentation ftp pit-manager.mit.edu pub/usenet/news.answers/gopherfaq

Gopher/Veronica documentation ftp

ftp.cso.uiuc.edu doc/net/uiucnet/vol6no1.txt

Gopher Newsgroup news comp.infosystems.gopher

Hyper-G

Definition: Hyper-G is a large-scale, multi-user distributed hypermedia information system

Action: U ::*(()):  G(L <= C <- S[F]..S[F])

Hyper-G via WWW http iicm.tu-graz.ac.at

ROOT 80

Hytelnet

Definition: Hytelnet provides Internet access info by hypertext (on network or TSR for DOS);

Action: Internet:  U ::  I(L <= C <- S[F]); DOS: U ::  I(L)

Hytelnet description ftp

ftp.usask.ca pub/hytelnet/README

Hytelnet via WWW http

www.cc.ukans.edu hytelnet html/START.TXT.html

Hytelnet via telnet telnet info.ccit.arizona.edu

‘hytelnet’

Jughead

Definition: Jughead gets menu information from gopher servers; Action: Jughead description ftp boombox.micro.umn.edu pub/gopher/Unix/GopherTools/jug

head/jughead.ReadMe

Knowbot

Definition: Knowbot provides a uniform interface to heterogeneous remote directory services;

Action: U ::  I(L <-<= S[F..F])

Knowbot via telnet telnet info.cnri.reston.va.us

185

Moo Gopher

Definition: Moo Gopher is a MUD-based gopher interface; Action: Moo Gopher blurb ftp mailbase.ac.uk pub/lists/unite/files/jays hous

e.txt

Moo Gopher via telnet telnet jh.ccs.neu.edu

1709

Mosaic for X/DOS/Mac

Definition: Mosiac is a client for network distributed hypermedia information and discovery;

Action: U :*(()):  G(L <= S[F]..S[F]); U := M <- S[F]; U ::  I(L <= S[F])

Mosaic home page http www.ncsa.uiuc.edu SDG/Software/Mosaic/NCSAMosaicH ome.html

Mosaic FAQ http www.ncsa.uiuc.edu SDG/Software/Mosaic/Docs/mosaic -faq.html

Mosaic ftp

ftp.ncsa.uiuc.edu Mosaic/README.Mosaic

Mosaic  unix xmosaic

Netfind

Definition: Netfind is a research prototype that provides a simple Internet ‘white pages’ user directory;

Action: U ::  I(L <= S[F])

Netfind description ftp

ftp.cs.colorado.edu pub/cs/distribs/netfind/README

Netfind via telnet telnet mudhoney.micro.umn.edu

‘netfind’

Netfind via email mail listserv@brownvm.brown.edu “get netfind help”

Netserv

Definition: Netserve is a server for access to data files and programs of general interest;

Action: U == M2 <- S[F] <- M1 <= U

Netserv mail netserv@frmop11.bitnet “get netserv helpfile”

Nslookup

Definition: Nslookup is an interactive program to query Internet domain name servers (gives IP address);

Action: U := M <= H

Nslookup man page unix

man nslookup

Nslookup command unix nslookup domain-name

Ping

Definition: Ping requests echo from network host; see if remote host is up;

Action: U := M <= H

Ping definition ftp

uxc.cso.uiuc.edu utils/ping/README

Ping via email mail

b.liddicott@ic.ac.uk Subject: #ping HOST@DOMAIN

Ping demo unix /usr/etc/ping -s uwm.edu

Prospero

Definition: Prospero provides user-centered view of remote files;

Action: U := F <+ S[F]

Prospero document ftp prospero.isi.edu pub/prospero/doc/prospero-user-

manual.tex.Z

Prospero documentation ftp prospero.isi.edu pub/prospero/doc/README-

prosper o-documents

Soft Pages

Definition: Soft Pages aids in retrieval of documents, software, and other resources from servers Action: Soft Pages documentation ftp ftp.tohoku.ac.jp

pub/spp

Project description ftp cs.ucl.ac.uk osi-ds/osi-ds-39-00.txt

Telnet

Definition: Telnet allows a user to login to a remote computer to use applications;

Action: U ::  I(L <= C <- H[A])

Telnet how ftp ftp.sura.net pub/nic/network.service.guides/

how.to.telnet.guide

TELNET ftp

nic.merit.edu documents/rfc/rfc0959.txt

Telnet demo telnet downwind.sprl.umich.edu

3000

Veronica

Definition: Veronica locates titles of Gopher items by keyword search;

Action: U ::  I(L <= C <- S[F]); U := M <- S[F]

Veronica definition ftp

veronica.scs.unr.edu veronica/veronica-faq

Gopher/Veronica documentation ftp

ftp.cso.uiuc.edu doc/net/uiucnet/vol6no1.txt

Veronica via gopher gopher

veronica.scs.unr.edu

Wide Area Information Server (WAIS)

Definition: WAIS retrieves resources by searching indexes of databases;

Action: U ::  I(L <= C <- S[F]); U := M <- S[F]

WAIS description ftp

quake.think.com wais/doc/overview.txt

WAIS document ftp

quake.think.com wais/doc/wais-corp.txt

WAIS documentation ftp quake.think.com

wais/doc/

WAIS via telnet telnet

quake.think.com

WAIS via telnet telnet wais.com

‘wais’

WAIS via Mosaic http www.ncsa.uiuc.edu SDG/Software/Mosaic/Interfaces/

wais/wais-interface.html

WAIS newsgroup news

comp.infosystems.wais

WAIS clients ftp sunsite.unc.edu pub/wais/clients/

WHOIS

Definition: WHOIS provides info on registered network names;

Action: U := M <= S[F]

Whois description ftp

gopher.ucdavis.edu pub/IETF/WNILS/Discussion.Paper

Whois specs ftp

nic.merit.edu documents/rfc/rfc0954.txt

Whois Servers List ftp

sipb.mit.edu pub/whois/whois-servers.list

Whois via email mail

service@nic.ddn.mil Subject: whois HOST.DOMAIN

Whois via telnet telnet

rs.internic.net

Whois via Unix unix whois ”string”

World Wide Web (WWW)

Definition: WWW retrieves resources by hyptertext browser of databases;

Action: U ::  I(L <= C <- S[F]); U := M <- S[F]

WWW definition http info.cern.ch hypertext/WWW/LineMode/Defaults

/default.html

WWW guide http

pulua.hcc.hawaii.edu guide/www.guide.html

WWW Bib http info.cern.ch hypertext/WWW/Bibliography.html 80

WWW faqs/guides http

cui www.unige.ch OSG/FAQ/www.html

WWW FAQ http

www.vuw.ac.nz non-local/gnat/www-faq.html

WWW Tools http info.cern.ch hypertext/WWW/Tools/Overview.ht ml

WWW via telnet telnet

info.cern.ch

WWW via email mail listserv@info.cern.ch

“www URL”

WWW Newsgroup news comp.infosystems.www

X.500

Definition: X.500 provides globally distributed directory service;

Action: U ::  I(L <= S[F])

X.500 Definition ftp

nic.merit.edu documents/rfc/rfc1308.txt

X.500 Implement ftp

nic.merit.edu documents/fyi/fyi 11.txt

X.500 via telnet telnet

rs.internic.net

X.500 via gopher gopher barbarian.rs.itd.umich.edu 7777

3 COMPUTER-MEDIATED COMMUNICATION (CMC) FORUMS

Email

Definition: Email allows a user to send message(s) to another user (or many users via mailing lists);

Action: R == M <- H[M <= U]

Email How ftp ftp.sura.net pub/nic/network.service.guides/

how.to.email.guide

SMTP ftp

nic.merit.edu documents/rfc/rfc0821.txt

Email 101 ftp

mrcnext.cso.uiuc.edu etext/etext93/email025.txt

Finding Email addresses ftp sunsite.unc.edu pub/docs/about-the-net/libsoft/

email address.txt

College Email addresses ftp

pit-manager.mit.edu pub/usenet/soc.college/

InterNetwork Mail ftp

csd4.csd.uwm.edu pub/internetwork-mail-guide

Email echo mail b.liddicott@ic.ac.uk

#echo

Email demo unix mail user@host.domain

Interactive Webbing (IW)

Definition: people create a common space for network distributed multimedia writing.

Action: U ::== G(M <= C <-> S[F])

Free For All http

south.ncsa.uiuc.edu Free.html 8241

BBS Home Page http

indian.irsa.jrc.it BBS/home.html

Internet Relay Chat (IRC)

Definition: IRC provides real-time, many-many text discussion divided into channels;

Action: U ::  I(M <= C <-> S[R..R])

IRC definition ftp

nic.merit.edu documents/rfc/rfc1459.txt

IRC FAQ ftp cs.bu.edu

irc/support/alt-irc-faq

IRC guide ftp dorm.rutgers.edu pub/Internet.documents/irc.basi

c.guide

IRC documentation ftp cs.bu.edu

irc/support/IRCprimer1.1.txt

IRC via telnet telnet sci.dixie.edu

6677

IRC via Unix unix irc

IRC newsgroup news

alt.irc

IRC ircii news

alt.irc.ircii

IRC lamers news

alt.irc.lamers

IRC recovery news

alt.irc.recovery

IRC undernet news alt.irc.undernet

Internet Talk Radio (ITR)

Definition: ITR is an audio multicast on the Internet;

Action: U ::(()) M <- H[S]

ITR introduction ftp

sunsite.unc.edu pub/talk-radio/ITRintro.readme

ITR FAQ mail

info@radio.com

Internet Talk Radio sites mail

sites@radio.com

ITR FAQ ftp rftm.mit.edu pub/usenet/news.answers/interne

t-talk-radio/

ITR Newsgroup news alt.internet.talk-radio

LISTSERV

Definition: LISTSERV is mailing-list server for group communication;

Action: U == M2 <- S[F] <- M1 <= U

LISTSERV definition mail

listserv@uacsc2.albany.edu “send listserv memo”

LISTSERV tips mail

listserv@bitnic.bitnet “get listserv tips”

LISTSERV searching mail listserv@ulkyvm.bitnet “get database search”

Mbone

Definition: Mbone is a live audio and video multicast virtual network on top of Internet;

Action: U ::*(()) M <-> H[S]

Mbone definition ftp

venera.isi.edu mbone/faq.txt

Multiple-User Dialogue (MUD)

Definition: MUDs are real-time interaction usually for social role-playing;

Action: U ::  I(M <= C <-> S[R..R])

Mud FAQ ftp

ftp.math.okstate.edu pub/muds/misc/mud-faq/

MUD Lists ftp caisr2.caisr.cwru.edu

pub/mud/

MUD WWW Page http math.okstate.edu

mud.html 8001

MUDWHO via WWW http

math.okstate.edu mudwho.html 8001

Mud Announcements news

rec.games.mud.announce

Mud Administration news

rec.games.mud.admin

Mud Miscellaneous news

rec.games.mud.misc

Mud Tiny news

rec.games.mud.tiny

Mud Diku news rec.games.mud.diku

Talk

Definition: Talk provides real-time interactive text with another user;

Action: U ::  I(M <-> H[R])

Talk man page unix

man talk

Talk demo unix talk user@host.domain

USENET

Definition: USENET provides asynchronous text discussion on many topics separated into newsgroups;

Action: U := I(M <- S[F] <- ..  <- S[F])

USENET description ftp rtfm.mit.edu pub/usenet/news.answers/what-is -

usenet/part1

USENET documentation ftp

wuarchive.wustl.edu doc/misc/acn/acn4-5.txt.Z

NNTP ftp

nic.merit.edu documents/rfc/rfc0977.txt

Beginners group news

news.announce.newusers

Usenet addicts news

alt.usenet.addict

Usenet humor discussion news

alt.humor.best-of-usenet.d

Usenet humor news

alt.humor.best-of-usenet

Usenet culture news

alt.culture.usenet

Usenet recovery news

alt.usenet.recovery

Usenet reader unix

rn newsgroup-name

Usenet post via email mail hierarchy-group-name@cs.utexas. edu Subject: Your Subject “Your Co ntents”

Usenet via WWW http info.cern.ch hypertext/DataSources/News/Grou

ps/Overview.html

Usenet via gopher finger

lesikar@tigger.stcloud.msus.edu

Usenet periodics finger

nichol@stavanger.sgp.slb.com

Usenet posters mail mail-server@rtfm.MIT.EDU “send usenet-addresses/FIRST L AST”

4 PROTOCOLS/STANDARDS

Protocols are the basis for operating tools and forums on the Internet. In this section, I summarize some sources of information about Internet and other protocols. Protocols associated with individual tools are listed with the tools above.

INTERNET PROTOCOLS

IP ftp

nic.merit.edu documents/rfc/rfc0791.txt

TCP ftp

nic.merit.edu documents/rfc/rfc0793.txt

OTHER STANDARDS/PROTOCOLS

Telecom stds ftp rftm.mit.edu pub/usenet/news.answers/standar ds-faq

ISO Protocols news

comp.protocols.iso

OSI ftp rftm.mit.edu pub/usenet/news.answers/osi-pro tocols

5 REFERENCES

OTHER STANDARDS/PROTOCOLS

Internet Tools EARN ftp ns.ripe.net

earn/earn-resource-tool-guide.t xt

Internet Tools NIR ftp mailbase.ac.uk pub/lists/nir/files/nir.status.

report

Internet Systems UNITE ftp mailbase.ac.uk pub/lists/unite/files/systems-l

ist.txt

Net Mgt Tools ftp nic.merit.edu documents/fyi/fyi 02.txt

6 ACTION NOTATION

NOTATION: In order to summarize the action of these network tools,
I've developed a notation to express the essential information/
communication action of each of the tools.
NOUNS
A Application/software
C Client
F File(s)
H Host/computer
I Interface
G Graphical User Interface
L List
M Message(s)
R Receiver
S Server
U User/sender
VERBS
:= has read access only
== has read and write access
:: interacts with user interface (synchronous)
(()) sound in interface
* video interface
<- copys (over network) from
<-> copys (over network) to and from
<+ linked from
<= created by
[] contains
() consists of
CONJUNCTIONS
; and
NOUN PHRASES
S[F] Server containing file(s)
M(L) Message consisting of a List
S[U] Server containing users(s) = other users using this server
H[U] Host containing users(s) = other users logged into this host
NOTATION EXAMPLES:
File Transfer Protocol: U == F <- H[F]
Narration: The user gets read/write access to a copy of a
file that was copied over the network from a
host containing that file.
Electronic mail: R == M <- H[M <= U]
Narration: The Receiver gets read/write access to a copy of
the Message that was copied over the network from
an (originating) Host on which resides the original
Message created by the User (Sender).

7 DESCRIPTION OF ITEMS

Item Descriptions

Internet-tools describes Internet tools project and other formats of the information Alex description NIR (Networked Information Retrieval) description of Alex Archie description archie - An Electronic Directory Service for the Internet, by Deutsch, Emtage,

and Heelan

Archie manual page reference manual page for archie Archie via telnet a telnet demo of archie Archie via WWW Hypertext Archie Servers Archie via Unix Unix command line for hosts with installed archie client, search for STRING Archie via email use archie via email Astra Help

the help file for Astra

Astra documentation

Astra User Guide

Biomix description BIOMIX current status, by Marcus Pattloch

Cello source source code

Compass description Internet Compass-your network guide, CERFnet offers new service for easier

Internetting

Essence description The distribution includes the following files and directories Essence document Essence – A Resource Discovery System Based on Semantic File Indexing Finger description Unix manual page for finger Finger demo

find out about user decemj@rpi.edu

Finger Protocol The Finger User Information Protocol, by D. Zimmerman Finger via WWW access finger information on USER via World Wide Web Finger via telnet access finger via telnet HOST=host on which USER is located, then type USER FingerInfo

a script to get information via finger, by Scott Yanoff

Who fingered? a script that tries to find out who fingered you (a general solution is not possible) Ftp how to

A guide to using ftp

FTP

File Transfer Protocol

Ftp via Unix

Unix command for ftp, once in, type help

Ftp via email

get files at anonymous ftp sites via email.

Ftp via telnet site at Univ of IA File Compression File compression, archiving, and text¡-¿binary formats, by David Lemson Gopher guide

Internet Gopher User’s Guide, ed. Paul Linder

Gopher demo

a session demonstrating gopher at Univ of Minnesota

Gopher via email use a gopher via email Gopher source description of gopher client sources Hyper-G via WWW Hyper-G gateway at Graz Univ. of Technology Hytelnet description Peter Scott’s Hypertext database of publicly accessible Internet sites Jughead description description of current status, by Rhett ‘Jonzy’ Jones Knowbot via telnet use a knowbot service via telnet; at the prompt, type ‘help’ Moo Gopher blurb a short description of Moo gopher Mosaic home page describes Mosaic, Internet information browser and World Wide Web client Mosaic

description of how to download and run NCSA Mosaic

Mosaic bring up xmosaic on system with client installed Netfind description a way to find someone on the Internet Netfind via telnet using telnet to find someone on the Internet Netfind via email use email to find someone on the Internet Netserv

a server for access to data files and programs of general interest

Nslookup man page Unix man page Nslookup command command to find out about comand-name Ping definition description of Ping and its use Ping demo

check to see if host uwm.edu is up and roundtrip time

Prospero document Prospero user’s manual, by Neuman and Augart Prospero documentation

description of prospero documentation available

Project description Soft Pages Project, by Johannsen and Mansfield Telnet how

a guide to using telnet

TELNET

Telnet Protocol

Telnet demo

check the weather

Veronica definition frequently asked questions and answers about veronica Veronica via gopher Search ALL of Gopherspace using veronica WAIS description an overview of WAIS WAIS clients

client to WAIS in different interfaces (X)

Whois description describes whois Whois specs

NICNAME/WHOIS

Whois Servers List List of Internet whois servers, by Matt Power Whois via telnet at the prompt, enter whois host-name Whois via Unix the Unix command for whois, where ‘string’ is the organization/person name

or Internet domain name

WWW definition home page for WWW WWW guide

Entering the WWW–A Guide to Cyberspace by Kevin Hughes

WWW Bib

Bibliography for the World Wide Web

WWW faqs/guides a list by Oscar Nierstrasz WWW FAQ

gnat’s World Wide Web Frequently Asked Questions and Answers Lists

WWW Tools

Tools for WWW providers

WWW via telnet using WWW via telnet WWW via email obtain a web file (e.g., http) via email; URL = Uniform resource locator X.500 Definition rfc specifying X.500 X.500 via telnet at the prompt, enter X500whois host-name Email How

Introduction to Sending and Receiving Electronic Mail

SMTP

Simple Mail Transfer Protocol

Email 101

describes how to use email as well as other Internet features, by John Goodwin

Finding Email addresses

how do I find someone’s email address?

College Email addresses

information on email addresses at colleges

InterNetwork Mail methods of sending mail from one network to another, by John Chew and Scott

Yanoff

Email echo sends email back to you Email demo

Unix (UCB) mail command

Free For All

an example of IW–allows user to place an entry in a list of WWW entries

BBS Home Page An experiment in interactive webbing IRC definition Internet Relay Chat Protocol, by Oikarinen and Reed IRC FAQ

frequently asked questions and answers list from newsgroup

IRC guide

Basic Guide to IRC

IRC documentation real-time, many-many text discussion divided into channels IRC via telnet using telnet to Undernet IRC via Unix

Unix command for hosts with an installed IRC client

ITR introduction Internet Talk Radio, by Carl Malamud ITR FAQ

Frequently Asked Questions on ITR (Usenet FAQ)

LISTSERV definition

the help file for listserv

Mbone definition the frequently asked questions and answers list about Mbone Mud FAQ

frequently asked questions and answers about MUDs

MUD Lists

The Totally Unofficial List of Internet Muds

MUD WWW Page Experimental MUD Page, information on Muds MUDWHO via WWW

MUDWHO Access via WWW

Talk man page Unix manual page for talk Talk demo

where user@host.domain is known

USENET description

What is Usenet?

USENET documentation

special issue of Amateur Computerist Newsletter about Usenet

NNTP

Network News Transfer Protocol

Beginners group for people new to USENET Usenet reader nn, xrn)mand for hosts with newsreader, rn (other readers Usenet post via email

post to newsgroup via email

Usenet via WWW access newsgroups via WWW Usenet periodics list of Usenet periodic postings Usenet posters find when and from what host person with name FIRST LAST posted to Usenet IP

Internet Protocol

TCP

Transmission Control Protocol

Telecom stds

International Standards for telecommunication

ISO Protocols

newsgroup

OSI

Open Systems Interconnection protocols

Internet Tools EARN

guide to network resource tools

Internet Tools NIR A status report on networked information retrieval tools and groups Net Mgt Tools Tools for Monitoring and Debugging TCP/IP Internets and Interconnected

Devices